how many liters of 75% gas and 90% gas would you need for 1200 liters of 85% gas?
Answer by checkley77(12844)

.90x+.75(1,200-x)=.85*1,200
.90x+900-.75x=1,020
.15x=1,020-900
.15x=120
x=120/.15
x=800 liters of 90% gas is used.
1,200-800=400 liters of 75% gas is used.
Proof:
.90*800+.75*400=1,020
720+300=1,020
1,020=1,020